# Weak-to-Strong Search: Align Large Language Models via Searching over Small Language Models

### v1

@@ Abstract
Large language models are usually fine-tuned to align with human preferences. However, fine-tuning a large language model can be challenging. In this work, we introduce weak-to-strong search, framing the alignment of a large language model as a test-time greedy search to maximize the log-likelihood difference between small tuned and untuned models while sampling from the frozen large model. This method serves both as (i) a compute-efficient model up-scaling strategy that avoids directly tuning the large model and as (ii) an instance of weak-to-strong generalization that enhances a strong model with weak test-time guidance. Empirically, we demonstrate the flexibility of weak-to-strong search across different tasks. In controlled-sentiment generation and summarization, we use tuned and untuned gpt2s to effectively improve the alignment of large models without additional training. Crucially, in a more difficult instruction-following benchmark, AlpacaEval 2.0, we show that reusing off-the-shelf small models (e.g., zephyr-7b-beta and its untuned version) can significantly improve the length-controlled win rates of both white-box and black-box large models against gpt-4-turbo (e.g., $34.4\rightarrow 37.9$ for Llama-3-70B-Instruct and $16.0\rightarrow 20.1$ for gpt-3.5-turbo-instruct), despite the small models’ low win rates $\approx 10.0$.
@@ 1 Introduction
Learning-based algorithms [1, 2, 3, 4, 5] have become the standard approach for aligning large language models (LLMs) with human preferences [3, 6, 7, 8, 9, 10]. However, fine-tuning large language models is resource-intensive and difficult to implement [4]. These challenges have motivated recent studies on search-based algorithms that keep the large language models frozen and steer their decoding with test-time guidance [11, 12, 13, 14, 15, 16, 17]. Typical examples of search-based algorithms include rejection sampling [16, 17] and Monte Carlo Tree Search [18, 19]. These search-based algorithms are promising as they can reuse the same learned guidance to steer the decoding of any large language model without additional training. However, existing search-based methods either simplify the search over tokens as a bandit problem [16, 17], which limits their steerability, or require a value function learned from scratch to address preference reward sparsity and reduce search depth [13, 18, 14], which can be as difficult as fine-tuning a large language model.
@@
To make search-based algorithms better suited for aligning large language models, we introduce weak-to-strong search, a simple algorithm that frames the alignment of a large model as a test-time search over the log-likelihoods of small language models. This algorithm makes two contributions: (1) First, it builds on the theoretical foundation of the token-level MDP for alignment [20], using the log-likelihood difference between small tuned and untuned language models as both reward and critic [4, 20] to guide the decoding of a large model (Section 4.1). Theoretically, this formulation is suitable for search as it converts the otherwise sparse preference reward to a per-token dense reward, which can be summed up as a value function (critic) [20]. Practically, this formulation allows the reuse of off-the-shelf small tuned and untuned language model pairs as steering forces, avoiding the need to train a reward or critic model from scratch. (2) Second, it introduces a beam search variant, Chunk-level Beam Search (CBS), tailored for optimizing the proposed search objective. CBS guides the large language model towards high-reward regions by alternating between sampling from the frozen large model and expanding promising states as evaluated by the small tuned and untuned models (Section 4.2). Especially, when the small models are weaker than the large model, our method can be viewed as an instance of weak-to-strong generalization [21] that makes the strong model stronger with weak test-time guidance (Figure 1).
@@
Figure 1: Weak-to-strong search enhances the alignment of large models with test-time guidance from small models (dashed lines). This method is applicable to white-box models that use the same or different vocabularies as the small models, as well as to black-box models. We present the results for the instruction-tuned models from each family (e.g., Llama2-7B denotes Llama-2-7b-chat).
@@
Empirically, we verify weak-to-strong search’s flexibility in various tasks (Section 5). First, in controlled-sentiment generation [22] and summarization [2], our method uses small language models of 124M parameters (i.e., gpt2) to effectively steer much larger language models from the GPT-2 (e.g., gpt2-xl[23], Llama-2 [7] and Llama-3 [24] families, at least as effective as existing methods. Then, in a more difficult instruction-following benchmark, AlpacaEval 2.0 [25], we show reusing off-the-shelf small models (e.g., zephyr-7b-beta and its untuned version) as test-time guidance can significantly improve the length-controlled win rates of both white-box and black-box large models against gpt-4-turbo (e.g, $34.4\rightarrow 37.9$ for Llama-3-70B-Instruct, and $16.0\rightarrow 20.1$ for gpt-3.5-turbo-instruct), despite the small models’ low win rates $\approx 10.0$ (Figure 1).
@@ 2 Related Work
Framing alignment as a test-time search to maximize a reward function is not a novel formulation. However, most existing works either simplify autoregressive decoding as a bandit problem [16, 17], which limits their steerability, or require a value function learned from scratch to handle sparse preference rewards and reduce search depth [13, 18], which can be as difficult as training a large language model from scratch. Our work avoids these issues by parametrizing the reward function with the log-likelihood difference between small tuned and untuned language models [4]. This parametrization not only simplifies the search objective, allowing a simple greedy search algorithm to generate good results, but also reuses off-the-shelf models as steering forces, eliminating the need to train a reward or critic model from scratch.
Concurrently with our work, Rafailov et al. [20] proposes a token-level MDP interpretation for language model alignment, demonstrating that a greedy likelihood search over a trained language model can achieve improvements over regular decoding. Our work builds on their theoretical foundations and proposes a novel greedy search algorithm designed for weak-to-strong guidance.
The idea of using small language models to align large language models has arisen in many recent works. The most related is proxy or emulated fine-tuning [34, 12, 11, 35], which uses the distributional difference of a small tuned and untuned model pair to modify the output distribution of a large model, approximating the output of the directly tuned large model. However, these methods require that both small and large models share the same vocabulary, limiting their practical applications. In contrast, our approach does not modify the sampling distribution of the large model at the token level. Instead, we perform a tree search that periodically selects the most promising states for further expansion (as evaluated by the small models) while sampling from the frozen large model’s distribution. Thus our approach does not require shared vocabulary and is applicable to black-box language models.
@@ 3.1 Aligning Language Models with Human Preferences
where $p({\mathbf{x}})$ is a distribution of prompts, ${\mathbf{y}}$ is the complete language model response, $r$ is a reward function that encourages human-preferred responses, and $\mathbb{D}_{\mathrm{KL}}$ limits how far the optimized language model $\pi$ can deviate from the reference (untuned) model $\pi_{\text{ref}}$. There are two main categories of alignment algorithms: (1) search-based algorithms that optimize Eq. 1 with graph-based search during inference  [16, 13, 18, 19, 11, 12], and (2) learning-based algorithms that optimize Eq. 1 through gradient descent, aiming for a parametrized optimal language model [1, 36, 4, 37]. Our work falls in the first category, proposing a search-based algorithm capable of using small language models to guide the decoding of a large language model to align with human preferences.
@@ 3.2 Duality between Language Models and Reward Functions
where $Z({\mathbf{x}})=\sum_{\mathbf{y}}{\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})}\exp\left(\frac{1}{\beta}r({\mathbf{x}},{\mathbf{y}})\right)$ denotes the partition function. One takeaway from this duality is that we can always express a reward function using tuned and untuned language models: (1) If a reward function is given [1, 2, 3], we can first obtain the optimally tuned language model under this reward function with any learning-based algorithms, and then use the tuned and untuned models $(\pi^{*},\pi_{\text{ref}})$ to reparametrize the reward function [42] (Eq. 3); (2) If a dataset is given from which the reward function can be derived, we can then directly parametrize the reward function with the tuned and untuned language models $(\pi^{*},\pi_{\text{ref}})$ during reward modeling [4] (Eq. 3).
@@ 4 Weak-to-Strong Search
In this section, we introduce weak-to-strong search, a search-based algorithm that aligns a large language model by searching over the log-likelihood difference between small tuned and untuned language models. First, we discuss how using language models to (re)parametrize the reward function from Eq. 1 makes the challenging objective solvable by a simple greedy search algorithm (e.g., beam search) (Section 4.1). Then, we introduce a practical beam search variant, called Chunk-level Beam Search (CBS) (Section 4.2), for optimizing the proposed objective, which is applicable to steering both white-box and black-box large language models.
4.1 Language Models as Both Reward and Critic
One practical challenge for search-based alignment algorithms is the sparsity of the preference reward signal. The preference reward function $r({\mathbf{x}},{\mathbf{y}})$, based on the Bradley-Terry model [43], only emits a terminal reward when the model response is complete. Search-based algorithms often struggle without any intermediate rewards or a critic model predicting future returns [44, 45]. However, if we parameterize this sparse reward function with language models (Section 3.2), we can obtain both a dense reward function and a critic function simultaneously.
Language models are a dense reward function.
To obtain a dense reward function, we leverage the duality between the sparse preference reward and the dense language model probability (Eq. 3). By explicitly factorizing the log-likelihood of a complete response ${\mathbf{y}}$ under the language models, we can obtain a sum-of-rewards style formulation for Eq. 3:
$r({\mathbf{x}},{\mathbf{y}})=\beta\left(\sum_{t=1}^{|{\mathbf{y}}|}\log\frac{\pi^{*}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}{\pi_{\text{ref}}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}\right)+\beta\log Z({\mathbf{x}}),$ (4)
where ${\mathbf{y}}_{<t}$ denotes the response tokens from $1$ to $t-1$, and the last response token $y_{|{\mathbf{y}}|}$ is always the EOS token. Combining Eq. 1 and 4, we can rewrite the original objective with a per-token reward:
$\operatorname*{arg\,max}_{\pi}\quad$ $\mathbb{E}_{{\mathbf{x}}\sim p({\mathbf{x}}),{\mathbf{y}}\sim\pi({\mathbf{y}}\mid{\mathbf{x}})}\left[\sum_{t=1}^{|{\mathbf{y}}|}\log\frac{\pi^{*}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}{\pi_{\text{ref}}(y_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}\right]$ (5a)
@@ 4.1 Language Models as Both Reward and Critic
where $\beta$ and $Z({\mathbf{x}})$ are omitted as they do not influence the optimal solution. It is important to note that the reference model that parametrizes the reward function ($\pi_{\mathrm{ref}}$) (Eq. 5a) and the reference model that constrains the test-time search space ($\pi_{\mathrm{base}}$) (Eq. 5b) can be different. Practically, decoupling the reference models is useful as it allows using a tuned and untuned language model pair - namely ($\bm{\pi^{*}}$, $\bm{\pi}_{\text{ref}}$) - to steer the decoding of any language model $\bm{\pi}_{\text{base}}$ without retraining.
@@
Partial return under language models is an implicit V-value function (critic) in the original KL-constrained sparse reward setting.
In Appendix A, we show that (inspired heavily by [20])
$\log\frac{\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}\propto\begin{cases}-V^{*}({\mathbf{x}})+V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime}),&\text{ if }y^{\prime}_{|{\mathbf{y}}^{\prime}|}\neq\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is incomplete)}\\ -V^{*}({\mathbf{x}})+r({\mathbf{x}},{\mathbf{y}}^{\prime}),&\text{ if }y^{\prime}_{|{\mathbf{y}}^{\prime}|}=\texttt{EOS}\text{ (${\mathbf{y}}^{\prime}$ is complete)}\end{cases}$ (6)
where $V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})$ denotes the value function, predicting the expected terminal reward under the optimal $\pi^{*}$ in the original KL-constrained sparse reward setting. Although $V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})$ is not necessarily achievable by the searched policy, it approximates how good the state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ is in the long run. In other words, continuing from the state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ of high partial return $\log\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})$ is likely to generate a complete response ${\mathbf{y}}$ with high overall return $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$.

### v3

@@ Abstract
Large language models are usually fine-tuned to align with human preferences. However, fine-tuning a large language model can be challenging. In this work, we introduce weak-to-strong search, framing the alignment of a large language model as a test-time greedy search to maximize the log-probability difference between small tuned and untuned models while sampling from the frozen large model. This method serves both as (1) a compute-efficient model up-scaling strategy that avoids directly tuning the large model and as (2) an instance of weak-to-strong generalization that enhances a strong model with weak test-time guidance. Empirically, we demonstrate the flexibility of weak-to-strong search across different tasks. In controlled-sentiment generation and summarization, we use tuned and untuned gpt2s to improve the alignment of large models without additional training. Crucially, in a more difficult instruction-following benchmark, AlpacaEval 2.0, we show that reusing off-the-shelf small models (e.g., zephyr-7b-beta and its untuned version) can improve the length-controlled win rates of both white-box and black-box large models against gpt-4-turbo (e.g., $34.4\%\rightarrow 37.9\%$ for Llama-3-70B-Instruct and $16.0\%\rightarrow 20.1\%$ for gpt-3.5-turbo-instruct), despite the small models’ low win rates $\approx 10.0\%$.
@@ 1 Introduction
Learning-based algorithms [1, 2, 3, 4, 5] have become the standard approach for aligning large language models (LLMs) with human preferences [3, 6, 7, 8, 9, 10]. However, fine-tuning large language models is resource-intensive and difficult to implement [4]. These challenges have motivated recent studies on search-based algorithms that keep the large language models frozen and steer their decoding with test-time guidance [11, 12, 13, 14, 15, 16, 17]. Typical examples of search-based algorithms include rejection sampling [16, 17] and Monte Carlo Tree Search [18, 19]. These search-based algorithms are promising as they can reuse the same guiding signal to steer the decoding of any large language model without additional training. However, existing search-based methods either simplify the search over tokens as a bandit problem [16, 17], which limits their steerability, or require a value function learned from scratch to address preference reward sparsity and prune search space [13, 18, 14], which can be as difficult as fine-tuning a large language model.
@@
To make search-based algorithms better suited for aligning large language models, we introduce weak-to-strong search, a simple algorithm that frames the alignment of a large model as a test-time search over the log-probabilities of small language models. This algorithm makes two contributions: (1) First, it builds on the theoretical foundation of the token-level MDP for alignment [20], using the log-probability difference between small tuned and untuned language models as both reward and value [4, 20] to guide the decoding of a large model (Section 4.1). Theoretically, this formulation is suitable for search as it converts the otherwise sequence-level sparse preference reward function to a per-token dense reward function, which can be summed up as a value function [20]. Practically, this formulation allows the reuse of off-the-shelf small tuned and untuned language model pairs as steering forces, avoiding the need to train a reward or value model from scratch. (2) Second, it introduces a beam search variant, Chunk-level Beam Search (CBS), tailored for optimizing the proposed search objective. CBS guides the large language model towards high-reward regions by alternating between sampling from the frozen large model and expanding promising states as evaluated by the small tuned and untuned models (Section 4.2). Especially, when the small models are weaker than the large model, our method can be viewed as an instance of weak-to-strong generalization [21] that makes the strong model stronger with weak test-time guidance (Figure 1).
@@
Figure 1: Weak-to-strong search enhances the alignment of large models through test-time guidance from small models (dashed lines). This method is applicable to white-box models that use the same or different vocabularies as the small models, as well as to black-box models. We present the results for the instruction-tuned models from each family (e.g., Llama2-7B denotes Llama-2-7b-chat).
@@
Empirically, we verify weak-to-strong search’s flexibility in various tasks (Section 5). First, in controlled-sentiment generation [22] and summarization [2], our method uses small language models of 124M parameters (i.e., gpt2) to effectively steer much larger language models from the GPT-2 (e.g., gpt2-xl[23], Llama-2 [7] and Llama-3 [24] families, at least as effective as existing methods. Then, in a more difficult instruction-following benchmark, AlpacaEval 2.0 [25], we show reusing off-the-shelf small models (e.g., zephyr-7b-beta and its untuned version) as test-time guidance can significantly improve the length-controlled win rates of both white-box and black-box large models against gpt-4-turbo (e.g, $34.4\%\rightarrow 37.9\%$ for Llama-3-70B-Instruct, and $16.0\%\rightarrow 20.1\%$ for gpt-3.5-turbo-instruct), despite the small models’ low win rates $\approx 10.0\%$ (Figure 1).
@@ 2 Related Work
Framing alignment as a test-time search to maximize a reward function is not a novel formulation. However, most existing works either simplify autoregressive decoding as a bandit problem [16, 17], which limits their steerability, or require a value function learned from scratch to handle sparse preference rewards and prune search space [13, 18], which can be as difficult as training a large language model from scratch. Our work avoids these issues by parametrizing the sparse preference reward function with the log-probability difference between small tuned and untuned language models [4]. This parametrization not only simplifies the search objective, allowing a simple greedy search algorithm to generate good results, but also reuses off-the-shelf models as steering forces, eliminating the need to train a reward or critic model from scratch.
Concurrently with our work, Rafailov et al. [20] proposes a token-level MDP interpretation for language model alignment, demonstrating that a greedy probability search over a trained language model can achieve improvements over regular decoding. Our work builds on their theoretical foundations and proposes a practical greedy search algorithm designed for weak-to-strong guidance.
The idea of using small language models to align large language models has arisen in many recent works. The most related is proxy or emulated fine-tuning [34, 12, 11, 35], which uses the distributional difference of a small tuned and untuned model pair to modify the output distribution of a large model, approximating the output of the directly tuned large model. However, these methods require that both small and large models share the same vocabulary, limiting their practical applications. In contrast, our approach does not modify the sampling distribution of the large model at the token level. Instead, we perform a tree search that periodically prioritizes the most promising states for further expansion (as evaluated by the small models) while sampling from the frozen large model’s distribution. Thus our approach does not require shared vocabulary and is applicable to black-box language models.
@@ 3.1 Aligning Language Models with Human Preferences
where $p({\mathbf{x}})$ is a distribution of prompts, ${\mathbf{y}}$ is the complete language model response, $r$ is a preference reward function that encourages human-preferred responses, and $\mathbb{D}_{\mathrm{KL}}$ limits how far the optimized language model $\pi$ can deviate from the reference (untuned) model $\pi_{\text{ref}}$. There are two main categories of alignment algorithms: (1) search-based algorithms that optimize Eq. 1 with graph-based search during inference  [16, 13, 18, 19, 11, 12], and (2) learning-based algorithms that optimize Eq. 1 through gradient descent, aiming for a parametrized optimal language model [1, 36, 4, 37]. Our work falls in the first category, proposing a search-based algorithm capable of using small language models to guide the decoding of a large language model to align with human preferences.
@@ 3.2 Duality between Language Models and Reward Functions
where $Z({\mathbf{x}})=\sum_{\mathbf{y}}{\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})}\exp\left(\frac{1}{\beta}r({\mathbf{x}},{\mathbf{y}})\right)$ denotes the partition function. One takeaway from this duality is that we can always express a reward function using tuned and untuned language models: (1) If a reward function is given [1, 2, 3], we can first obtain the optimally tuned language model under this reward function with any learning-based algorithms, and then use the tuned and untuned models $(\pi^{*},\pi_{\text{ref}})$ to reparametrize the reward function [42]; (2) If a dataset is given from which the reward function can be derived, we can then directly parametrize the reward function with the tuned and untuned language models $(\pi^{*},\pi_{\text{ref}})$ during reward modeling [4].
@@ 4 Weak-to-Strong Search
In this section, we introduce weak-to-strong search, a search-based algorithm that aligns a large language model by searching over the log-probability difference between small tuned and untuned language models. First, we discuss how using language models to parametrize the preference reward function (Eq. 1) makes the reward-maximization problem solvable by a simple greedy search algorithm (e.g., beam search) (Section 4.1). Then, we introduce a practical beam search method, Chunk-level Beam Search (CBS) (Section 4.2), that balances reward maximization and KL minimization, which is applicable to steering both white-box and black-box large language models.
4.1 Language Models as Both Reward and Value Functions
One practical challenge for search-based alignment algorithms is the sparsity of the preference reward signal. The preference reward function $r({\mathbf{x}},{\mathbf{y}})$, based on the Bradley-Terry model [43], only emits a terminal reward when the model response is complete. Search-based algorithms often struggle without any intermediate rewards or a value function providing intermediate guidance [44, 45]. However, if we parameterize this sparse reward function with language models (Section 3.2), we can obtain both a dense reward function and a value function simultaneously.
Language models as a dense reward function.
To obtain a dense reward function, we leverage the duality between the sparse preference reward function and the dense language model probability (Eq. 3). By explicitly factorizing the log-probability of a complete response ${\mathbf{y}}$ under the language models, we obtain a sum-of-rewards style formulation for Eq. 3:
$r({\mathbf{x}},{\mathbf{y}})=\beta\left(\sum_{t=1}^{|{\mathbf{y}}|}\log\frac{\pi^{*}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}{\pi_{\text{ref}}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}\right)+\beta\log Z({\mathbf{x}}),$ (4)
where ${\mathbf{y}}_{<t}$ denotes the response tokens from $1$ to $t-1$, and the last response token ${\textnormal{y}}_{|{\mathbf{y}}|}$ is the EOS token. Combining Eq. 1 and 4, we rewrite the original objective with a per-token reward function:
$\operatorname*{arg\,max}_{\pi}\quad$ $\mathbb{E}_{{\mathbf{x}}\sim p({\mathbf{x}}),{\mathbf{y}}\sim\pi({\mathbf{y}}\mid{\mathbf{x}})}\left[\sum_{t=1}^{|{\mathbf{y}}|}\log\frac{\pi^{*}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}{\pi_{\text{ref}}({\textnormal{y}}_{t}\mid{\mathbf{x}},{\mathbf{y}}_{<t})}\right]$ (5a)
@@ 4.1 Language Models as Both Reward and Value Functions
where $\beta$ and $Z({\mathbf{x}})$ are omitted as they do not influence the optimal solution. It is important to note that the reference model that parametrizes the reward function ($\pi_{\mathrm{ref}}$) (Eq. 5a) and the reference model that constrains the test-time search space ($\pi_{\mathrm{base}}$) (Eq. 5b) can be different. Practically, decoupling the reference models is useful as it allows using a tuned and untuned language model pair - namely ($\bm{\pi^{*}}$, $\bm{\pi}_{\text{ref}}$) - to steer the decoding of any base language model $\bm{\pi}_{\text{base}}$ without retraining.
@@
Cumulative reward under language models as a value function [20].
Appendix A shows that:
$\log\frac{\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}{\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})}\propto\begin{cases}-V^{*}({\mathbf{x}})+V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})&\text{ if ${\mathbf{y}}^{\prime}$ is incomplete}\\ -V^{*}({\mathbf{x}})+r({\mathbf{x}},{\mathbf{y}}^{\prime})&\text{ if ${\mathbf{y}}^{\prime}$ is complete},\end{cases}$ (6)
where $V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})$ denotes the value function, predicting the expected terminal reward under the optimal $\pi^{*}$ in the original KL-constrained sparse reward problem. Although $V^{*}({\mathbf{x}},{\mathbf{y}}^{\prime})$ is not necessarily achievable by the searched policy, it approximates how good the state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ is in the long run. In other words, continuing from the state $({\mathbf{x}},{\mathbf{y}}^{\prime})$ of high cumulative reward $\log\pi^{*}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}^{\prime}\mid{\mathbf{x}})$ is likely to generate a complete response ${\mathbf{y}}$ with high overall return $\log\pi^{*}({\mathbf{y}}\mid{\mathbf{x}})-\log\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}})$.